OpenStudy (anonymous):

Find the difference: 4/5 − 2/7

OpenStudy (anonymous):

Change to the common fraction of in this case 35/ so 4/5 = 28/35 and 2/7 =10/35 28/35 - 10/35 = 18/35

TheSmartOne (thesmartone):

Wrong subject, please post in the mathematics subject.